# FlowMotion: Target-Predictive Flow Matching for Realistic Text-Driven Human Motion Generation

### v1

@@ Abstract
Achieving highly diverse and perceptually consistent 3D character animations with natural motion and low computational costs remains a challenge in computer animation. Existing methods often struggle to provide the nuanced complexity of human movement, resulting in perceptual inconsistencies and motion artifacts. To tackle these issues, we introduce FlowMotion, a novel approach that leverages Conditional Flow Matching (CFM) for improved motion synthesis. FlowMotion incorporates an innovative training objective that more accurately predicts target motion, reducing the inherent jitter associated with CFM while enhancing stability, realism, and computational efficiency in generating animations. This direct prediction approach enhances the perceptual quality of animations by reducing erratic motion and aligning the training more closely with the dynamic characteristics of human movement. Our experimental results demonstrate that FlowMotion achieves higher balance between motion smoothness and generalization capability while maintaining the computational efficiency inherent in flow matching compared to state-of-the-art methods.
@@ 1 Introduction
The synthesis of 3D human body motion has diverse applications across fields such as robotics [1, 2, 3], VR/AR [4, 5], entertainment [6, 7, 8, 9], and social interaction in virtual 3D spaces [10, 11, 12, 13]. To achieve realistic and context-aware motion synthesis, motion generation techniques frequently leverage data-driven motion capture data. These techniques include, for instance, approaches based on deep learning [14, 15], reinforcement learning [16, 17], and hybrid methods [18], which utilize these datasets to model and predict complex movements.
@@
A significant subset of techniques for synthesizing 3D human body motion has emerged, driven by recent advances in generative models. These methods have efficiently enabled the generation of motion conditioned on user-defined inputs, with a primary focus on descriptive texts, that specifies the intended movement. Such approaches leverage the adaptability of generative models to produce contextually appropriate and plausible motion sequences  [19, 20, 21, 22, 23, 24]. While recent advances in 3D human motion generation are significant, several challenges remain. The inherent complexities of motion generation are exacerbated by the need to incorporate diverse constraints, such as spatial trajectories [25, 26], interactions with surrounding objects [27, 28], or temporal specifications defined by keyframes [29], all aimed at producing lifelike movements.
@@
Among these generative methods, classes such as variational autoencoders (VAEs) [21, 22, 30], generative adversarial networks (GANs) [31, 23, 32, 24], and diffusion models [20, 19, 33] have received great attention for 3D human motion generation, each offering distinct advantages and limitations. While VAEs have been used in motion generation, their output diversity is often limited due to the issue of posterior collapse. Similarly, GANs can suffer from mode collapse, making it challenging to achieve sufficient diversity in the generated motions. In contrast, diffusion models excel in generating diverse and natural motions, albeit with increased latency during the generation process.
@@
To address these limitations, hybrid models have emerged to leverage the strengths of different architectures. Recent proposals include integrating features from GANs and diffusion models to enhance generation speed without sacrificing generalization capability [34]. Furthermore, there are also approaches combining the strengths of GANs, VAEs, and diffusion models, utilizing distillation techniques for even faster sampling [35]. Additionally, diffusion techniques have been applied to latent spaces to enable faster sampling, using VAEs [36].
One of the main advantages of generative methods is their capability of producing a diverse range of plausible human motion sequences, enabling users to explore multiple interpretations of a desired motion and select the sequence that best aligns with their creative intent. This flexibility is particularly valuable in scenarios requiring user-controlled motion generation.
@@
Beyond the aforementioned generative approaches, flow-matching-based models have demonstrated promising performance in generating both images [37, 38] and 3D human motion [39], achieving fast synthesis with adequate diversity in the results. However, in our analysis of the flow matching approach by Hu et al. [39] for 3D human body motion synthesis, we observed temporal inconsistencies (jitter effect) [4], which can compromise smoothness and perceptual fidelity. We hypothesize that this artifact arises from the conditional flow matching (CFM) formulation [37], where the calculation of the learned vector field inherently incorporates variations resulting from the differences between the data representations and samples from a standard Gaussian distribution used during training.
To take advantage of the computational efficiency of flow matching while avoiding the jitter effect, we propose a novel training objective explicitly designed to approximate the original, unperturbed motion, effectively attenuating fluctuations within the vector field during synthesis. Additionally, we employ the acceleration metric to rigorously quantify the temporal smoothness of the synthesized motion and ensure a high degree of fidelity in the generated results.
Our approach presents the following main contributions:
  • 1.
    Target-oriented training objective: We introduce a novel, target-oriented training objective for CFM that directly predicts the target motion. This approach outperforms diffusion and flow matching methods in temporal coherence and fidelity.
  • 2.
    Temporal smoothness and jitter reduction: By introducing a novel training objective and leveraging the acceleration metric, we ensure smooth transitions and minimize undesirable tremors in the generated motion.
  • 3.
    Balance between fidelity and smoothness: Our approach achieves a superior trade-off between generation fidelity, evaluated using Fréchet Inception Distance (FID), and temporal smoothness, assessed by jitter metrics (Fig. 1), resulting in reliable and natural motion sequences.
@@
This paper is organized as follows: In Sec. 2, we review the related work on generative methods for 3D human motion synthesis. In Sec. 3, we present our method, covering both theoretical foundations and computational details. Sec. 4 describes the experiments conducted to evaluate the effectiveness of our approach. In Sec. 5, we discuss and analyze the results of these experiments. Finally, we conclude our study in Sec. 6, highlighting potential directions for future work based on the insights gained from our approach.
Figure 1: FID-Jitter evaluation on the HumanML3D dataset. FlowMotion achieves a favorable balance between motion smoothness, as quantified by lower Jitter, and generation fidelity, as measured by lower Fréchet Inception Distance (FID) scores, demonstrating its ability to generate high-quality motion compared to prior methods.
@@ 2 Related Work
More recently, Guo et al. [30] introduced a temporal VAE enhanced with attention mechanisms for generating human motion from text, alongside the widely adopted HumanML3D dataset and associated evaluation metrics. This dataset, one of the largest currently available, has become a standard benchmark for subsequent research, including the present work.
Subsequent advancements have seen the incorporation of Transformer architectures. ACTOR [41] utilized a Transformer-based VAE, while TEACH [42] conditioned motion generation on sequences of textual motion descriptions using a similar architectural approach. Further, TM2T [43] established connections between text tokens and motion sequences via Transformers, demonstrating efficacy on both HumanML3D and KIT-ML [44] datasets. T2M-GPT [21] employed a Vector Quantized Variational AutoEncoder (VQ-VAE) in conjunction with a CLIP text encoder [45] to achieve a discrete representation of movement, exhibiting robust generalization capabilities. Text2Gestures [46] utilizes a Transformer encoder to process descriptive text encoded with GloVe [47] and a Transformer decoder which, combined with previous movements, generates subsequent ones, enabling the production of expressive animations.
@@ 2.2 Diffusion and Flow Matching Models
However, these diffusion-based approaches are often slow in the inference process due to the extensive sequence of steps required for sampling. Efforts like MLD [36] improve inference speed by performing diffusion in a latent space achieved through a prior VAE process, reducing computational resources and accelerating the inference process. Similarly, MotionLCM [35] and EMDM [34] achieve real-time inference by accelerating the sampling process, skipping denoising steps in the diffusion process through distillation and a discriminator process similar to those used in GANs, respectively. To ensure coherent body generation and prevent artifacts such as levitation, PhysDiff [18] integrate physical constraints during the sampling phase of the diffusion framework, though such constraints may reduce generation speed.
On the other hand, text-conditioned human motion generation via flow matching has shown promising results [39]. Using the same architecture as MDM, this approach achieves better generalization and faster inference due to the streamlined trajectory in the sampling process. However, the use of flow matching relies on intractable integrals. Thus, in practice, one uses the Conditional Flow Matching (CFM) model [37], but it tends to generate erratic or jittering motions during inference. The variations introduced by the CFM formulation, where the learned vector field is influenced by the interpolation between data representations and random Gaussian noise, may contribute to this artifact. In our method, we propose reducing jittering by modifying the loss function of the CFM, such that it directly compares the generated image with the noise-free original image.
Distinct from previous work, our results indicate that the technique here presented delivers visually superior results without increasing computational cost.
@@ 3 The FlowMotion Method
Our method generates realistic human motion sequences conditioned on textual descriptions, with focus on producing smooth and controlled movements. To achieve this, we leverage Conditional Flow Matching (CFM) [37], which enables the learning of complex motion distributions. In contrast to standard flow matching approaches, CFM operates on conditional distributions, which is advantageous for sampling as it promotes straighter trajectories and therefore, faster sampling.
Our training objective directly predicts the target motion, enhancing stability and reducing jitter compared to prior CFM and diffusion-based approaches. By employing a Transformer-based architecture, our method efficiently processes motion and text embeddings, enabling the generation of motion sequences that exhibit strong alignment with the provided textual descriptions.
In this section, we provide a concise overview of flow matching, followed by the details of our proposed CFM-based framework, including the training procedure and sampling strategy.
@@ 3.1 Flow Matching
Given a finite set of samples $x_{1}$ from an unknown data distribution $q$, and initializing with $p_{0}=\mathcal{N}(0,I)$, Lipman et al. [37] introduce the concept of a conditional probability path $p_{t}(\cdot|x_{1}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$, defined for each sample $x_{1}$. At the final time $t=1$, $p_{1}(\cdot|x_{1})$ is defined as a Gaussian distribution centered at $x_{1}$ with a small standard deviation $\sigma_{\min}\geq 0$, concentrating the probability around the sample. The distribution $p_{t}(x)$ is then defined as the marginalization of these conditional probability paths:
@@
The original Conditional Flow Matching (CFM) objective [37] is formulated with respect to the initial sample $x_{0}$. However, for our iterative sampling process, we propose redefining the objective in terms of the intermediate state $x_{t}$. Utilizing the conditional vector field construction from Lipman et al. [37], which expresses the vector field based on the means, standard deviations, and their derivatives. Thus, we adopt their formulation to define the conditional vector field:
@@
Rather than training a model to approximate $u_{t}$ directly, we propose predicting the target $x_{1}$ through a neural network conditioned on $x_{t}$ (Sec. 3.2.1). Given the predicted $x_{1}$ and the fixed $\sigma_{\text{min}}$, the vector field $u_{t}$ is computed via Eq. (4). This approach bypasses explicit vector field estimation while preserving the theoretical guarantees of CFM.
@@ 3.2 Framework
While our model implements the architecture of MDM [20], it is crucial to highlight that the input $x_{t}$ is derived from a linear interpolation process, following the CFM framework (Eq. (3)), rather than the stochastic corruption process characteristic of diffusion models such as MDM.
@@ 3.2.1 Training
Following a similar approach proposed by Ramesh et al. [49] for diffusion models, where they modified the loss function to predict the target $x_{1}$ directly—achieving improved performance compared to predicting the noise—we tailor this concept to our CFM framework. Thus, our training objective is defined as:
@@
Figure 3: Overview of the training process. In each epoch, the process starts with a sample $x_{1}$ from the training dataset and a sample $x_{0}\sim\mathcal{N}(0,{I})$. An intermediate representation $x_{t}$ is then determined via linear interpolation between $x_{1}$ and $x_{0}$. The region highlighted in red denotes the space of valid human motions. Note that at the beginning of the process, both $x_{0}$ and $x_{t}$ can be found outside this space.
@@ 3.3 Model
Methods RP Top3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MMModality $\uparrow$ Jitter $\rightarrow$
Real motion $0.797^{\pm.001}$ $0.001^{\pm.000}$ $2.973^{\pm.005}$ $9.488^{\pm.081}$ $47.262$^{\pm$0.069$}$
TM2T [43] $0.726^{\pm.002}$ $1.507^{\pm.019}$ $3.469^{\pm.011}$ $8.526^{\pm.086}$ $2.540^{\pm.063}$ $87.208$^{\pm$0.151$}$
T2M-GPT [21] $0.776^{\pm.005}$ $0.124^{\pm.008}$ $3.131^{\pm.019}$ $9.567^{\pm.101}$ $1.719^{\pm.120}$ $76.046$^{\pm$1.123$}$
MDM [20] $0.706^{\pm.004}$ $0.519^{\pm.050}$ $3.658^{\pm.022}$ $9.442^{\pm.070}$ $2.871^{\pm.047}$ $46.30^{\pm.9}$
MotionGPT [60] $0.659^{\pm.003}$ $0.185^{\pm.007}$ $4.013^{\pm.022}$ $9.291^{\pm.067}$ $3.481^{\pm.119}$ $94.770$^{\pm$0.250$}$
MotionLCM [35] $0.802^{\pm.002}$ $0.416^{\pm.010}$ $3.008^{\pm.005}$ $9.740^{\pm.069}$ $2.123^{\pm.076}$ $37.578$^{\pm$0.224$}$
MLD [36] $0.758^{\pm.003}$ $0.427^{\pm.012}$ $3.269^{\pm.015}$ $9.775^{\pm.073}$ $2.573^{\pm.089}$ $22.772$^{\pm$0.113$}$
MFM [39] $0.666^{\pm.005}$ $0.446^{\pm.047}$ $5.186^{\pm.023}$ $9.987^{\pm.089}$ $2.419^{\pm.096}$ $576.679$^{\pm$3.153$}$
FlowMotion-Big (ours) $0.648^{\pm.004}$ $0.268^{\pm.028}$ $5.319^{\pm.032}$ $9.796^{\pm.069}$ $2.327^{\pm.077}$ $40.81^{\pm.5}$
FlowMotion (ours) $0.663^{\pm.005}$ $0.278^{\pm.030}$ $5.239^{\pm.027}$ $9.788^{\pm.103}$ $2.349^{\pm.055}$ $39.522$^{\pm$0.770$}$

### v2

# FlowMotion: Target-Predictive Conditional Flow Matching for Jitter-Reduced Text-Driven Human Motion Generation

@@ Abstract
Achieving high-fidelity and temporally smooth 3D human motion generation remains a challenge, particularly within resource-constrained environments. We introduce FlowMotion, a novel method leveraging Conditional Flow Matching (CFM). FlowMotion incorporates a training objective within CFM that focuses on more accurately predicting target motion in 3D human motion generation, resulting in enhanced generation fidelity and temporal smoothness while maintaining the fast synthesis times characteristic of flow-matching-based methods. FlowMotion achieves state-of-the-art jitter performance, achieving the best jitter in the KIT dataset and the second-best jitter in the HumanML3D dataset, and a competitive FID value in both datasets. This combination provides robust and natural motion sequences, offering a promising equilibrium between generation quality and temporal naturalness.
@@ 1 Introduction
The synthesis of 3D human body motion has diverse applications across fields such as robotics [1, 2, 3], VR/AR [4, 5], entertainment [6, 7, 8, 9], and social interaction in virtual 3D spaces [10, 11, 12, 13]. While recent advances in 3D human motion generation are significant, several challenges remain. The inherent complexities of motion generation are exacerbated by the need to incorporate diverse constraints, such as spatial trajectories [14, 15], interactions with surrounding objects [16, 17], or temporal specifications defined by keyframes [18], all aimed at producing lifelike movements.
To achieve realistic and context-aware motion synthesis, motion generation techniques frequently leverage data-driven motion capture data. These techniques include, for instance, approaches based on deep learning [19, 20], reinforcement learning [21, 22], and hybrid methods [23], which utilize these datasets to model and predict complex movements.
@@
Recent studies in generative models have driven the development of new techniques for synthesizing 3D human body motion. These methods primarily focus on generating realistic movements based on user inputs, especially descriptive text that specifies the intended action. A key advantage of these generative approaches is their ability to produce a diverse range of plausible motion sequences from a single prompt. This allows users to explore multiple interpretations of a desired movement and select the sequence that best aligns with their creative vision. Among these generative methods, classes such as variational autoencoders (VAEs) [24, 25, 26], generative adversarial networks (GANs) [27, 28, 29, 30], and diffusion models [31, 32, 33] have received great attention for 3D human motion generation, each offering distinct advantages and limitations. While VAEs have been used in motion generation, their output diversity is often limited due to the issue of posterior collapse. Similarly, GANs can suffer from mode collapse, making it challenging to achieve sufficient diversity in the generated motions. In contrast, diffusion models excel in generating diverse and natural motions, albeit with increased latency during the generation process.
@@
Hybrid models, on the other hand, aim to minimize the limitations of individual architectures by combining their complementary strengths. Recent proposals include integrating features from GANs and diffusion models to enhance generation speed without sacrificing generalization capability [34]. Furthermore, there are also approaches combining the strengths of GANs, VAEs, and diffusion models, utilizing distillation techniques for even faster sampling [35]. Additionally, diffusion techniques have been applied to latent spaces to enable faster sampling, using VAEs [36].
@@
Beyond the aforementioned generative approaches, flow-matching-based models have demonstrated promising performance in generating both images [37, 38] and 3D human motion [39], achieving fast synthesis with adequate diversity in the results. While the motion synthesis approach by Hu et al. [39] performs well when adequate computational resources (specifically GPU memory enabling larger batch sizes) are available, we observed challenges under hardware constraints. The necessity of using smaller batch sizes due to limited GPU memory can produce temporal inconsistencies, resulting in jitter effect in the generated motion. Even with careful tuning, achieving stable training with the standard conditional flow matching formulation presented difficulties under these conditions. We hypothesize that this artifact, potentially linked to the inherent variations in the conditional flow matching (CFM) formulation [37] itself (due to differences between complex data representations and the target Gaussian distribution), is amplified by the training instability introduced when using smaller batch sizes necessitated by hardware limitations.
To harness the computational efficiency of flow matching while addressing the jitter effect, we introduce a training objective designed to approximate the original, unperturbed motion—effectively reducing fluctuations in the vector field during synthesis. This objective is implemented within the Conditional Flow Matching framework and focuses on directly predicting the target motion, leading to a more stable and accurate vector field. This approach is conceptually aligned with a recent formulation introduced by Lipman et al. [40]. However, our novel contribution lies in the application of this method to 3D human motion generation, specifically tailored to mitigate jitter.
To quantitatively evaluate the temporal smoothness of the generated motion, we employ an acceleration-based metric, which offers an objective and interpretable measure of motion stability and fidelity.
Figure 1: FID-Jitter evaluation on the HumanML3D and KIT datasets. Jitter represents the absolute difference from the ground truth Jitter. HumanML3D is represented in light blue, KIT in orange. Stars of the same color denote our results: FlowMotion. FlowMotion proximity to the Ground Truth (GT), indicated in yellow at the origin, highlights its effectiveness in achieving a balance between generation fidelity, measured by Fréchet Inception Distance (FID), and motion smoothness, quantified by lower Jitter values, indicating a generation closer to real-world motion compared to existing methods.
Our work introduces a novel application of a CFM training objective that directly predicts target motion to 3D human motion generation, offering enhanced generation fidelity and temporal smoothness. While maintaining a lightweight model architecture suitable for resource-constrained environments, our approach also retains the fast synthesis times characteristic of flow-matching-based methods. The application of this training objective to 3D human motion generation results in a method, which we call FlowMotion. Experimental results indicate that FlowMotion achieves a promising equilibrium between generation quality (FID) and temporal naturalness (Jitter) (Fig. 1). FlowMotion exhibits top-tier performance, achieving the best jitter in the KIT dataset and the second-best jitter in the HumanML3D dataset, while attaining a very competitive FID value in both datasets. This closely approximates real human motion, bolstering the ability to generate smooth, realistic animations. This enhanced fidelity and smoothness highlight the potential of our approach to generate robust and natural motion sequences.
@@
This paper is organized as follows: In Sec. 2, we present the related work on generative methods for 3D human motion synthesis. In Sec. 3, we present our method, covering both theoretical foundations and computational details. Sec. 4 describes the experiments conducted to evaluate the effectiveness of our approach. In Sec. 5, we discuss and analyze the results of these experiments. Finally, we conclude our study in Sec. 6, highlighting potential directions for future work based on the insights gained from our approach.
@@ 2 Related Work
More recently, Guo et al. [26] introduced a temporal VAE enhanced with attention mechanisms for generating human motion from text, along with the widely adopted HumanML3D dataset and associated evaluation metrics. This dataset, in conjunction with the KIT-ML dataset [42], represents one of the largest currently available and has become a standard benchmark for subsequent research, including the present work.
Subsequent advancements have seen the incorporation of Transformer architectures. ACTOR [43] utilized a Transformer-based VAE, while TEACH [44] conditioned motion generation on sequences of textual motion descriptions using a similar architectural approach. Further, TM2T [45] established connections between text tokens and motion sequences via Transformers, demonstrating efficacy on both HumanML3D and KIT-ML [42] datasets. T2M-GPT [24] employed a Vector Quantized Variational AutoEncoder (VQ-VAE) in conjunction with a CLIP text encoder [46] to achieve a discrete representation of movement, exhibiting robust generalization capabilities. Similarly, MotionGPT [47], based on the same VQ-VAE, constructs a vocabulary of movements combined with an advanced Transformer (T5 [48]) for integration with text. Text2Gestures [49] utilizes a Transformer encoder to process descriptive text encoded with GloVe [50] and a Transformer decoder which, combined with previous movements, generates subsequent ones, enabling the production of expressive animations.
@@ 2.2 Diffusion and Flow Matching Models
However, these diffusion-based approaches tends to be slow in the inference process due to the extensive sequence of steps required for sampling. Efforts like MLD [36] improve inference speed by performing diffusion in a latent space achieved through a prior VAE, reducing computational resources and accelerating inference. Similarly, MotionLCM [35] and EMDM [34] achieve real-time inference by accelerating sampling, omitting denoising steps in the diffusion pipeline through distillation and a discriminator, respectively, in a manner analogous to techniques used in GANs. To ensure coherent body generation and prevent artifacts such as levitation, PhysDiff [23] integrate physical constraints during the sampling phase of the diffusion framework, though such constraints may reduce generation speed.
On the other hand, text-conditioned human motion generation via flow matching has shown promising results [39]. Using the same architecture as MDM, this approach achieves better generalization and faster inference due to the streamlined trajectory in the sampling process. However, the use of flow matching relies on intractable integrals. Thus, in practice, one uses the Conditional Flow Matching (CFM) model [37], which can generate erratic or jittering motions during inference, particularly under limited computational resources. The variations introduced by the CFM formulation, where the learned vector field is influenced by the interpolation between data representations and random Gaussian noise, may contribute to this artifact. In our method, we propose reducing jitter by modifying the loss function of the CFM, such that it directly compares the generated image with the noise-free original image.
Distinct from previous work, our results indicate that the technique presented herein delivers visually superior results, exhibiting both higher generation fidelity and improved temporal smoothness, without increasing computational cost.
@@ 3 The FlowMotion Method
Our method, FlowMotion, focuses on generating text-conditioned human motion sequences with high fidelity and temporal smoothness. To achieve this equilibrium between generation quality and naturalness, we leverage Conditional Flow Matching (CFM) [37], which enables the learning of complex motion distributions. In contrast to standard flow matching approaches, CFM operates on conditional distributions, which is advantageous for sampling as it promotes straighter trajectories and therefore, faster sampling.
Our training objective directly predicts the target motion, enhancing stability and reducing jitter compared to prior CFM and diffusion-based approaches. By employing a Transformer-based architecture, our method efficiently processes motion and text embeddings, enabling generating motion sequences exhibiting strong diversity with the provided textual descriptions.
In this section, we provide a concise overview of flow matching, followed by the details of our CFM-based framework, including the training procedure and sampling strategy.
@@ 3.1 Flow Matching
Given a finite set of samples $x_{1}$ from an unknown data distribution $q$, and initializing with $p_{0}=\mathcal{N}(0,I)$, Lipman et al. [37] introduce the concept of a conditional probability path $p_{t}(\cdot|x_{1}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$, defined for each sample $x_{1}$. At the final time $t=1$, $p_{1}(\cdot|x_{1})$ is defined as a Gaussian distribution centered at $x_{1}$ with a small standard deviation $\sigma_{\min}>0$, concentrating the probability around the sample. The distribution $p_{t}(x)$ is then defined as the marginalization of these conditional probability paths:
@@
The original Conditional Flow Matching (CFM) objective [37] is formulated with respect to the initial sample $x_{0}$. However, for our iterative sampling process, we redefine the objective in terms of the intermediate state $x_{t}$. Utilizing the conditional vector field construction from Lipman et al. [37], which expresses the vector field based on the means, standard deviations, and their derivatives. Thus, we adopt their formulation to define the conditional vector field:
@@
Rather than training a model to approximate $u_{t}$ directly, we predict the target $x_{1}$ through a neural network conditioned on $x_{t}$ (Sec. 3.2.1). A similar formulation has also been recently proposed by Lipman [40] as an alternative parameterization of the vector field $u_{t}$. Given the predicted $x_{1}$ and the fixed $\sigma_{\text{min}}$, the vector field $u_{t}$ is computed via Eq. (4). This approach bypasses explicit vector field estimation while preserving the theoretical guarantees of CFM.
@@ 3.2 Framework
While our model implements the architecture of MDM [31], it is important to highlight that the input $x_{t}$ is derived from a linear interpolation process, following the CFM framework (Eq. (3)), rather than the stochastic corruption process characteristic of diffusion models such as MDM.
@@ 3.2.1 Training
Building upon the findings of Ramesh et al. [52], which demonstrated that directly predicting the target $x_{1}$ improved performance over noise prediction—a strategy subsequently adopted by the diffusion-based MDM—we adapt this approach to our Conditional Flow Matching framework. Employing a similar network architecture as MDM, our training objective is thus defined as:
@@
Figure 3: Overview of the training process. In each epoch, the process starts with a sample $x_{1}$ from the training dataset and a sample $x_{0}\sim\mathcal{N}(0,{I})$. An intermediate representation $x_{t}$ is then determined via linear interpolation between $x_{1}$ and $x_{0}$. The region highlighted in red denotes the space of valid human motions. Note that at the beginning of the process, $x_{t}$ can be found outside this space.
@@ 3.3 Model
Methods #params $\downarrow$ FID $\downarrow$ Diversity $\uparrow$ Jitter $\rightarrow$ FIDJ $\downarrow$ MM-Dist $\downarrow$ MMModality $\uparrow$ RP Top3 $\uparrow$
Real motion - $0.001^{\pm.000}$ $9.488^{\pm.081}$ $47.262$^{\pm$0.069$}$ $0.000$ $2.973^{\pm.005}$ $0.797^{\pm.001}$
TM2T [45] 43.0M $1.507^{\pm.019}$ $8.526^{\pm.086}$ $87.208$^{\pm$0.151$}$ $16.542$ $3.469^{\pm.011}$ $2.540^{\pm.063}$ $0.726^{\pm.002}$
MDM [31] 17.9M $0.519^{\pm.050}$ $9.442^{\pm.070}$ $46.30^{\pm.9}$ $5.084$ $3.658^{\pm.022}$ $2.871^{\pm.047}$ $0.706^{\pm.004}$
MFM [39] 17.9M $0.446^{\pm.047}$ $9.987^{\pm.089}$ $576.679$^{\pm$3.153$}$ $32.906$ $5.186^{\pm.023}$ $2.419^{\pm.096}$ $0.666^{\pm.005}$
MLD [36] 26.4M $0.427^{\pm.012}$ $9.775^{\pm.073}$ $22.772$^{\pm$0.113$}$ $3.391$ $3.269^{\pm.015}$ $2.573^{\pm.089}$ $0.758^{\pm.003}$
MotionLCM [35] 39.7M $0.416^{\pm.010}$ $9.740^{\pm.069}$ $37.578$^{\pm$0.224$}$ $3.742$ $3.008^{\pm.005}$ $2.123^{\pm.076}$ $0.802^{\pm.002}$
MotionGPT [47] 332.8M $0.185^{\pm.007}$ $9.291^{\pm.067}$ $94.770$^{\pm$0.250$}$ $4.148$ $4.013^{\pm.022}$ $3.481^{\pm.119}$ $0.659^{\pm.003}$
T2M-GPT [24] 228.4M $0.124^{\pm.008}$ $9.567^{\pm.101}$ $76.046$^{\pm$1.123$}$ $2.615$ $3.131^{\pm.019}$ $1.719^{\pm.120}$ $0.776^{\pm.005}$
FlowMotion-Big 26.3M $0.268^{\pm.028}$ $9.796^{\pm.069}$ $40.81^{\pm.5}$ $2.400$ $5.319^{\pm.032}$ $2.327^{\pm.077}$ $0.648^{\pm.004}$
FlowMotion 17.9M $0.278^{\pm.030}$ $9.788^{\pm.103}$ $39.522$^{\pm$0.770$}$ $2.453$ $5.239^{\pm.027}$ $2.349^{\pm.055}$ $0.663^{\pm.005}$